CH/DOC/94/08

ANU Running

Plans for July 1994

Author: B.R. Fulton

Date: 26 April 1994

Plans for July running at ANU

This provides a summary of the (still evolving) plans for
running at ANU in July.  We still do not have firm dates for the
running periods and there may be a bit of movement.  I will
alert people as soon as we have actual dates.
For advance notice, it looks like the next running period after this
will be around February next year.  At this time experiments will
probably be needed for Glen (depending on current analysis), Becky
(depending on results of July tests) and Steve C (depending on current
analysis).  The February period matches the timescale for knowing what
will be needed for the above, and will also hopefully coincide with
installation work on the new beamline (if we can get it shipped out in
time).
In addition this period straddles the start of the next beamtime allocation,
so we can combine any remaining time from this year with some new time
(assuming we get some approved - PUBLICATIONS!!!!!!)

The running this July will include:
1. Work on DA system  Install new system software
Install and test FETCH module
Change to EUROGAM tape write system
Test new AIM modules
2. Experiment on 16O+16O to populate the 24Mg 12C+12C breakup
MF coordinate. Two hybrids. 4 days running. 12C and 16O beams
3. Experiment on 12C on 24Mg to see the 24Mg 12C+12C backangle
WNC coordinates. Two hybrids. 2 days running. 12C beam

4. Tests for proposed gamma measurements
NMC coordinates 4 NaI's. 3 days. 12C and 24Mg beams
5. Tests on MEGHA module with new Hammamatsu detectors installed
DLW coordinates. Commission then look for 24Mg 12C+12C breakup
4 days beam. 12C, 16O and 24Mg beams

Contact Roger if you need a visa
Person named as coordinator for the individual runs will be responsible
 for any specialist equipment needed plus targets.
Brian to check detector stocks.
Steve to ensure enough tapes at ANU

The present plan is to ship stuff out from Birmingham by 10 June (Roger
 will coordinate this with Dick, Vic and Steve providing the DA kit to
 go and Nobby the gamma detector test kit).
Vic and Dick will aim to be at ANU by 20 June to spend that week working
 on the DA system work.  The tests of the AIM modules will actually be
 done as part of the first experiments, with a bunch of 8-16 of the
 strip detector signals running through the normal system in parallel
 with the same signals through the AIM's and into the current TDC's
 (subject to channel capacity).
Brian, Martin, Nobby, Becky, Garry out to do setup in week beginning
 27 June.  (Nobby/Garry may delay arrival to stay later - see below)
Bill out for period approx 4-14 July
6 days for 16O+16O and 12C+24Mg experiments in week beginning 4 July
After this complete chamber changed over for MEGHA tests
Roger Doug and Paul arrive sometime around 18 July
4 days for gamma measurements in week beginning 18 July
3 days for MEGHA tests in week beginning 25 July

Brian will be away at Brisbane Conference 3-8 July.
Martin to return soon after 18 July
Brian, Garry, Nobby to return arround 22 July (Garry/Nobby may go out
 later and stay longer to help with MEGHA tests).
Brian Fulton
26.4.94